OpenStudy (hba):

If the area of a square is two more than twice the length of its diagonal then the length of one side of the square is........

hartnn (hartnn):

where are u stuck ? what u tried ?

OpenStudy (hba):

I got the answer root(2)-2 ?

hartnn (hartnn):

oh let me calculate and check.....

OpenStudy (hba):

Ok thanks :)

hartnn (hartnn):

is this your equation : s^2-2root2 s -2 = 0

OpenStudy (hba):

S^2-2roor2s+2=0

OpenStudy (hba):

root*

hartnn (hartnn):

i get root 2 +2

hartnn (hartnn):

how did u get -2 ?

OpenStudy (hba):

I just really dont know this gap year has made my mind rusty :( Let me check again and tell you :)

hartnn (hartnn):

just get the equation right S^2-2roor2s+2=0 is incorrect

hartnn (hartnn):

s^2-2root2 s -2 = 0

OpenStudy (hba):

Thanks Alot. :)

hartnn (hartnn):

oh, welcome. but did u get your error ?

OpenStudy (hba):

Well a little explanation would help ?

hartnn (hartnn):

sure, area of a square is two more than twice the length of its diagonal means if we add 2 to 'twice the length of its diagonal' then we can equate area of square and 'twice the length of its diagonal' so we get s^2 = 2root 2 s \(\huge +2\)

hartnn (hartnn):

got that ? ^

OpenStudy (hba):

Thank you very much :)

hartnn (hartnn):

welcome ^_^